# Local stability guarantees for data-driven quadratically nonlinear models

### v3

@@ Abstract
The Navier Stokes equations (NSEs) are partial differential equations (PDEs) to describe the nonlinear convective motion of fluids and they are computationally expensive to simulate because of their high nonlinearity and variables being fully coupled. Reduced-order models (ROMs) are simpler models for evolving the flows by capturing only the dominant behaviors of a system and can be used to design controllers for high-dimensional systems. However it is challenging to guarantee the stability of these models either globally or locally. Ensuring the stability of ROMs can improve the interpretability of the behavior of the dynamics and help develop effective system control strategies. For quadratically nonlinear systems that represent many fluid flows, the Schlegel and Noack trapping theorem [1] can be used to check if ROMs are globally stable (long-term bounded). This theorem was subsequently incorporated into system identification techniques that determine models directly from data [2]. In this work, we relax the quadratically energy-preserving constraints in this theorem, and then promote local stability in data-driven models of quadratically nonlinear dynamics. First, we prove a theorem outlining sufficient conditions to ensure local stability in linear-quadratic systems and provide an estimate of the stability radius. Second, we incorporate this theorem into system identification methods and produce a-priori locally stable data-driven models. Several examples are presented to demonstrate the effectiveness and accuracy of the proposed algorithm.
@@ I Introduction
In this work, we develop a theorem in the context of projection-based ROMs that enables us to promote local stability of data-driven models with quadratic nonlinearities. The theorem provides sufficient condition under which linear-quadratic systems with a relaxed quadratically energy-preserving structure are locally stable. We can also reproduce the Schlegel and Noack theorem [1] by tightening the relaxation of the energy-preserving constraints. Formalizing stability under weak breaking of the energy-preserving nonlinearity is critical because many flows exhibit boundary conditions such as inflow/outflow conditions. Providing conditions under which this class of systems is guaranteed to be stable is helpful for designing fluid flow control.
We next present how to promote the local stability in data driven models with a modified optimization loss function that is designed under the guidance of this theorem. The sparse identification of nonlinear dynamics (SINDy) [28] provides a number of readily available tools and is a useful platform for illustrating this approach.
This paper is organized as follows: Sec. II introduces the energy-preserving structure in the quadratic nonlinearities, followed by the Schlegel and Noack trapping theorem which connects the stability of the system with such structure. We then present a new theorem which derives an analytical estimate for the stability domain under the situation where the energy-preserving constraints are relaxed. In Sec. III, we define our “extended trapping SINDy” algorithm. The implementation is available through the PySINDy software package which is open-source [29, 30]. In Sec. IV, we demonstrate the efficacy and utility of this new technique for system identification and modeling on a number of examples. In Sec. V, we conclude our contribution and main achievement of this work with suggestions of future work.
@@ 1. $\bullet$
    bounded if there exists a constant $\rho>0$ that is independent of the initial time 00, and for every $c\in(0,\rho)$, there is $\mu=\mu(c)>0$ such that
@@ 4. $\bullet$
Let $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuously differentiable function. Then we can define the derivative of $V(x)$ along trajectories with respect to time [33, 34],
@@ 1. $\bullet$
    a Lyapunov function for the value of $c\in\mathbb{R}$ if $\dot{V}(\bm{x})\leq 0$ at all points of $V(\bm{x})=c$,
@@ 3. $\bullet$
An attracting trapping region is one in which trajectories in $\mathcal{D}\setminus\mathcal{N}$ eventually fall into $\mathcal{N}$, while a monotonically attracting trapping region is one in which trajectories in $\mathcal{D}\setminus\mathcal{N}$ are always falling towards $\mathcal{N}$. Fig. 1 illustrates this difference. These definitions form the basis for our analysis following in the next a few sections.
Figure 1: Sketch of an attracting trapping region. $\bm{\Gamma}_{1}$ (in red, solid) and $\bm{\Gamma}_{2}$ (in blue, solid) are two trajectories starting outside an attracting trapping region $\bm{\mathcal{N}}$ where all trajectories will eventually enter in and stay there forever. $V$ is a Lyapunov function and each dashed closed surface $V=V(\bm{x})=c_{i}\ (i=1,2,3,4)$ is a Lyapunov surface with $c_{1}>c_{2}>c_{3}>c_{4}$. For each Lyapunov surface outside $\bm{\mathcal{N}}$, $\bm{\Gamma}_{2}$ only passes through at most once, while $\bm{\Gamma}_{1}$ crosses $V=c_{3}$ more than once before each enters into $\mathcal{\bm{N}}$. If we consider the Lyapunov function as a form of energy, along $\bm{\Gamma}_{2}$, the energy is always decreasing before entering $\bm{\mathcal{N}}$. If all trajectories starting outside $\bm{\mathcal{N}}$ behave the same as $\bm{\Gamma}_{2}$, then $\bm{\mathcal{N}}$ is a monotonically attracting trapping region, otherwise an attracting trapping region.
@@ II.2 Flows with energy-preserving nonlinearities
$E_{i}$, $L_{ij}$, and $Q_{ijk}$ are tensors of time-independent coefficients, obtained from spatial inner products between the $\bm{\chi}_{i}(\bm{x})$ and the spatiotemporal operators that define the model dynamics. We have assumed that the pressure gradient term in Eq. (1) has been removed as in Rowley et al[35]. Moreover, Eq. (9) can be proved that it is locally Lipschitz since it is continuously differentiable over any compact domain in a Euclidean space; therefore the existence and uniqueness of the initial value problem are guaranteed. The class of systems considered for the original trapping theorem are those with energy-preserving nonlinearity, which is defined as
@@ II.3 Schlegel and Noack trapping theorem
Consider an effectively quadratically nonlinear system. There exists a monotonically trapping ball $\bm{B}(\bm{m},R_{m})$ if and only if the real, symmetric matrix $\bm{A}^{S}$ is negative definite (Hurwitz) with eigenvalues $\lambda_{r}\leq\cdots\leq\lambda_{1}<0$; the radius is then given by $R_{m}=\|\bm{d}\|_{2}/|\lambda_{1}|$. [^1] If a system is long-term bounded (not necessarily exhibiting a monotonically trapping region) and effectively nonlinear, the existence of an $\bm{m}$ producing a negative semidefinite $\bm{A}^{S}$ is still guaranteed.
@@ II.4 Analytical estimates of the local stability of linear-quadratic systems
The time derivative of the Lyapunov function given by Eq. (II.4) is
@@
We can then introduce $\tilde{Q}^{\prime}_{ijk}=\tilde{Q}_{ijk}+\tilde{Q}_{jki}+\tilde{Q}_{kij}$, and $\bm{H}_{0}=\mathcal{T}(\tilde{\bm{Q}}^{\prime})$. See that
@@ II.5 Interpretation of the local stability radius
The $\rho_{+}$ and $\rho_{-}$ are two nontrivial roots of Eq. (31c). As shown in the Fig. 2, $\rho_{+}$ is a conservative estimate of the diameter of the inscribed ball $\textbf{B}_{\rho_{+}}$ in the locally stable region $\bm{\Omega}_{\rho_{+}}$. Meanwhile $\rho_{-}$ is an estimate of the diameter of the set $\Omega_{\rho_{-}}$ and $\dot{v}=0$ at the boundary $\partial\Omega_{\rho_{-}}$. Negative energy growth in $\textbf{B}_{\rho_{+}}\setminus\textbf{B}_{\rho_{-}}$ which is an elliptical annulus in the state space of $\bm{y}$, is guaranteed according to Thm. 2 and inside $\textbf{B}_{\rho_{-}}$ there exists a region $\bm{\Omega}_{\rho_{-}}$, inside of which positive energy growth can be observed. Then the long-time behavior of solutions to the system (23) with initial conditions $\bm{P}^{\frac{1}{2}}\bm{y}_{0}=\bm{z}_{0}\in\textbf{B}_{\rho_{+}}$ can alternate between $\textbf{B}_{\rho_{-}}\setminus\bm{\Omega}_{\rho_{-}}$ and $\bm{\Omega}_{\rho_{-}}$, as shown in the Fig. 3. In a special case where $\bm{z}_{0}$ is a real solution to Eq. (31a), the long-time behavior of the system stays at the boundary of $\bm{\Omega}_{\rho_{-}}$. It is not difficult to conclude that
@@
which correctly reproduces the original trapping theorem guaranteeing global boundedness (in fact, a monotonically trapping region). We can also validate the limit by setting $\epsilon_{Q}$ to zero so that ~$\bm{\tilde{}}{\bm{H}}=0$, then the time derivative of the Lyapunov function (31a) becomes
@@
Figure 2: Sketch of the trapping region. $\dot{v}<0$ is observed in $\bm{\Omega}_{\rho_{+}}\setminus\bm{\Omega}_{\rho_{-}}$. $\partial\bm{\Omega}_{\rho_{+}}$ has a complex geometry but can be bounded below by $\partial\bm{B}_{\rho_{+}}$. $\bm{\Omega}_{\rho_{-}}$ has $\dot{v}>0$, a shape close to ellipsoid, and can be bounded above by $\partial\bm{B}_{\rho_{-}}$. Therefore long-term stability is guaranteed for any initial conditions $\|\bm{y}(t_{0})\|_{\bm{P}}=\|\bm{z}(t_{0})\|_{2}\leq\rho_{+}$.
@@
Note that the “shell” of stability $\bm{B}_{\rho_{+}}\setminus\bm{B}_{\rho_{-}}$ vanishes when $\rho_{+}=\rho_{-}$ or equivalently when the square root vanishes:
@@ III Extended trapping SINDy algorithm
In this section, we present how to incorporate Thm. 2 into system identification methods. System identification is the process of directly constructing dynamical models from data via optimization, rather than intrusive sampling of the fields and field derivatives as in traditional POD-Galerkin methods [32, 11].
In particular, we use the trapping SINDy algorithm [2] and formulate a modified “extended” trapping SINDy algorithm. We first review the standard SINDy algorithm [28] and the trapping SINDy algorithm, which successfully incorporates the trapping theorem (Thm. 1) into model identification in Sec. III.1. We then introduce our proposed extended trapping SINDy algorithm in Sec III.2.
@@ III.1 Standard and trapping SINDy algorithm
where $\dot{\bm{X}}_{:,j}$ denotes the $j^{th}$ column of $\dot{\bm{X}}$, $\|\cdot\|_{0}$ counts the number of nonzero elements and $\gamma$ determines the strength of sparsity-promotion during identification of the $\bm{\Xi}$. One may also notice that $\|\cdot\|_{0}$ is not convex, which leads to a nonconvex optimization. One way to ameliorate this is to instead use the $\ell_{1}$ norm,resulting in the LASSO [39]. The $\ell_{1}$ is convex but promotes a weaker form of sparsity. A computationally efficient alternative to directly solving the $\ell_{0}$-regularized problem was demonstrated in Brunton et al[28], where a sequential threshold least-squares algorithm was used to greedily solve the problem; its convergence properties were proven in [40].
To encode the enforced energy-preserving constraint on the quadratics terms, Loiseau et al[41] vectorizes $\bm{\Xi}[:]=\bm{\xi}\in\mathbb{R}^{rN}$ to recast the constrained minimization problem as an unconstrained convex problem using an augmented functional formulation. If the constraint is encoded as $\bm{C}\bm{\xi}=\bm{q}$, together with the improvements above, the optimization problem reads,
@@
where $\bm{\hat{\Theta}}(\bm{X})=\bm{I}\otimes\bm{\Theta}(\bm{X})$, $\dot{\bm{X}}[:]$ is the vectorized $\dot{\bm{X}}$ so that we have $\bm{\hat{\Theta}}\in\mathbb{R}^{rM\times rN}$, $\dot{\bm{X}}[:]\in\mathbb{R}^{rM}$. The number of constraints of the energy-preserving structure is given by $p=r(r+1)(r+2)/6$ and therefore we have $\bm{C}\in\mathbb{R}^{p\times rN}$, $\bm{q}\in\mathbb{R}^{p}$ and the Lagrange multiplier $\bm{\zeta}\in\mathbb{R}^{p}$. Additionally, for the linear equality constraints of the energy-preserving structure, $\bm{q}=0$ and the number of free parameters is given by $rN-p=2p$. If needed, these constraints can be used to impose any additional linear relationships from physical knowledge, such as enforcing global invariant preservation or enforcing a priori known values of any entries $\xi_{i}$.
@@
$\bm{\xi},\bm{\zeta},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}$ $\max_{\bm{\zeta}}\Biggr{[}\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (37)
@@
where the new term denotes that the largest eigenvalue of $\tilde{\bm{A}}^{S}$ will be enforced to be negative so that $\tilde{\bm{A}}^{S}$ is Hurwitz, and the magnitude of whose loss is modulated by the hyperparameter $\eta$. One should notice that the last term of this problem is not convex but convex composite, making the problem non-convex, however the details of the algorithm used for solving the optimization problem in (37) will not be introduced in this work; see Kaptanoglu et al[2] for more details.

### v4

# Extending the trapping theorem to provide local stability guarantees for quadratically nonlinear models

@@ Abstract
The Navier Stokes equations (NSEs) are partial differential equations (PDEs) to describe the nonlinear convective motion of fluids and they are computationally expensive to simulate because of their high nonlinearity and variables being fully coupled. Reduced-order models (ROMs) are simpler models for evolving the flows by capturing only the dominant behaviors of a system and can be used to design controllers for high-dimensional systems. However it is challenging to guarantee the stability of these models either globally or locally. Ensuring the stability of ROMs can improve the interpretability of the behavior of the dynamics and help develop effective system control strategies. For quadratically nonlinear systems that represent many fluid flows, the Schlegel and Noack trapping theorem [1] can be used to check if ROMs are globally stable (long-term bounded). This theorem was subsequently incorporated into system identification techniques that determine models directly from data [2]. While the Schlegel and Noack trapping theorem provides global stability criteria for systems with strictly energy-preserving nonlinearities, many physical systems, including those with inflow/outflow boundary conditions, exhibit weakly relaxed energy-preserving structures. This work introduces two key advances: (1) a theorem establishing analytical stability bounds for linear-quadratic systems under relaxed energy-preserving constraints, explicitly quantifying the local stability radius, and (2) the extended trapping SINDy algorithm, which embeds these theoretical guarantees into data-driven system identification. By integrating Lyapunov’s direct method with the trapping theorem framework, our approach enables the first provably locally stable models for quadratic dynamics with weakly broken energy-preserving nonlinearities. Several examples are presented to demonstrate the effectiveness and accuracy of the proposed algorithm.
@@ I Introduction
I.1 Contributions of this work
This work establishes a theorem within the framework of projection-based reduced-order models (ROMs), which guarantees local stability for data-driven models with quadratic nonlinearities. The theorem derives sufficient conditions for the local stability of linear-quadratic systems under weakly relaxed quadratic energy-preserving structures. Furthermore, we demonstrate that the Schlegel and Noack theorem [1] is recovered as a limiting case by restricting the relaxation of energy-preserving constraints. Formalizing stability under weakly violated energy-preserving nonlinearities is critical for applications involving boundary conditions such as inflow/outflow configurations, which are ubiquitous in fluid systems. Rigorous stability guarantees for this class of systems provide a foundation for designing robust fluid flow control strategies.
We next demonstrate how local stability in data-driven models can be enforced through a modified optimization loss function derived from the theorem. The sparse identification of nonlinear dynamics (SINDy) framework [28], known for its utility in parsimonious model discovery, serves as a natural platform for this implementation. We modify the original optimizer based on SINDy for identifying globally stable models [2] with minimal adjustments, adapting it to discover locally stable systems exhibiting weakly quadratic energy-preserving structures while enabling estimation of the local stability radius. Furthermore, we extend the framework by generalizing quadratic energy-preserving terms through the introduction of a Lyapunov matrix, thereby integrating Lyapunov’s direct method into the Schlegel and Noack trapping theorem.
The rest of the paper is organized as follows: Sec. II begins by reviewing energy-preserving quadratic nonlinearities and the Schlegel and Noack trapping theorem, which establishes stability criteria for systems with such structures. In Sec. II, we introduce a novel theorem that rigorously quantifies the stability domain for systems where energy-preserving constraints are relaxed—a critical advancement for flows with non-conservative boundary conditions (e.g., open flows). In Sec. III, we develop the extended trapping SINDy algorithm, a significant extension of the original trapping SINDy framework that augments it with local stability guarantees by embedding our theoretical advances into the optimization process. This algorithm is implemented in the open-source PySINDy package [29, 30], enabling reproducible and accessible stable model discovery. Sec. IV showcases the algorithm’s ability to recover stable models from noisy data, even under weakened energy-preserving assumptions, across several benchmark systems. Finally, Sec. V synthesizes the core innovations of this work: (1) the first analytical stability bounds for quadratically nonlinear systems with weakly energy-preserving structure, and (2) a generalizable framework for enforcing stability in data-driven models, bridging Lyapunov theory with the trapping theorem. We conclude with open challenges and future extensions.
@@ 1. $\bullet$
    bounded if there exists a constant $\rho>0$ that is independent of the initial time 0, and for every $c\in(0,\rho)$, there is $\mu=\mu(c)>0$ such that
@@ 4. $\bullet$
Let $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuously differentiable function. Then we can define the derivative of $V(\bm{x})$ along trajectories with respect to time [33, 34],
@@ 1. $\bullet$
    a Lyapunov function for the value of $c\in\mathbb{R}_{>0}$ if $\dot{V}(\bm{x})\leq 0$ at all points of $V(\bm{x})=c$,
@@ 3. $\bullet$
An attracting trapping region is one in which trajectories in $\mathcal{D}\setminus\mathcal{N}$ eventually fall into $\mathcal{N}$, while a monotonically attracting trapping region is one in which trajectories in $\mathcal{D}\setminus\mathcal{N}$ are always falling towards $\mathcal{N}$. Also, a trapping region is not exactly an attractor, as a trapping region is a neighborhood of an attractor $\mathcal{A}$ that satisfies $\phi_{t}(\mathcal{A})\subset\mathcal{A}$ for all $t$ and no proper subset of $\mathcal{A}$ is so. Fig. 1 illustrates this difference. Consider $V$ a Lyapunov function and each dashed closed surface $V=V(\bm{x})=c_{i}\ (i=1,2,3,4)$ is a Lyapunov surface with $c_{1}>c_{2}>c_{3}>c_{4}$. For each Lyapunov surface outside $\bm{\mathcal{N}}$, $\bm{\Gamma}_{2}$ only passes through at most once, while $\bm{\Gamma}_{1}$ crosses $V=c_{3}$ more than once before each enters into $\bm{\mathcal{N}}$. If we consider the Lyapunov function as a form of energy, along $\bm{\Gamma}_{2}$, the energy is always decreasing before entering $\bm{\mathcal{N}}$. If all trajectories starting outside $\bm{\mathcal{N}}$ behave the same as $\bm{\Gamma}_{2}$, then $\bm{\mathcal{N}}$ is a monotonically attracting trapping region, otherwise an attracting trapping region. These definitions form the basis for our analysis following in the next a few sections.
Figure 1: Sketch of an attracting trapping region. $\bm{\Gamma}_{1}$ (in red, solid) and $\bm{\Gamma}_{2}$ (in blue, solid) are two trajectories starting outside an attracting trapping region $\bm{\mathcal{N}}$ where all trajectories will eventually enter in and stay there forever.
@@ II.2 Flows with energy-preserving nonlinearities
$E_{i}$, $L_{ij}$, and $Q_{ijk}$ are tensors of time-independent coefficients, obtained from spatial inner products between the $\bm{\chi}_{i}(\bm{x})$ and the spatiotemporal operators that define the model dynamics. We have assumed that the pressure gradient term in Eq. (1) has been removed as in Rowley et al. [35]. Moreover, Eq. (9) can be proved that it is locally Lipschitz since it is continuously differentiable over any compact domain in a Euclidean space; therefore the existence and uniqueness of the initial value problem are guaranteed. The class of systems considered for the original trapping theorem are those with energy-preserving nonlinearity, which is defined as
@@ II.3 Schlegel and Noack trapping theorem
Consider an effectively quadratically nonlinear system with energy-preserving nonlinearity. There exists a monotonically trapping ball $\bm{B}(\bm{m},R_{m})$ if and only if the real, symmetric matrix $\bm{A}^{S}$ is negative definite (Hurwitz) with eigenvalues $\lambda_{r}\leq\cdots\leq\lambda_{1}<0$; the radius is then given by $R_{m}=\|\bm{d}\|_{2}/|\lambda_{1}|$. [^1] If a system is long-term bounded (not necessarily exhibiting a monotonically trapping region) and effectively nonlinear, the existence of an $\bm{m}$ producing a negative semidefinite $\bm{A}^{S}$ is still guaranteed.
@@ II.4 Analytical estimates of the local stability of linear-quadratic systems
The time derivative of the Lyapunov function given by Eq. (25) is
@@
[…]
@@ II.5 Interpretation of the local stability radius
The $\rho_{+}$ and $\rho_{-}$ are two nontrivial roots of Eq. (31c). As shown in the Fig. 2, $\rho_{+}$ is a conservative estimate of the diameter of the inscribed ball $\textbf{B}_{\rho_{+}}$ in the locally stable region $\bm{\Omega}_{\rho_{+}}$. $\partial\bm{\Omega}_{\rho_{+}}$ has a complex geometry but can be bounded below by $\partial\bm{B}_{\rho_{+}}$. Meanwhile $\rho_{-}$ is an estimate of the diameter of the set $\Omega_{\rho_{-}}$ and $\dot{v}=0$ at the boundary $\partial\Omega_{\rho_{-}}$. $\bm{\Omega}_{\rho_{-}}$ has $\dot{v}>0$, a shape close to ellipsoid, and can be bounded above by $\partial\bm{B}_{\rho_{-}}$. Negative energy growth in $\textbf{B}_{\rho_{+}}\setminus\textbf{B}_{\rho_{-}}$ which is an elliptical annulus in the state space of $\bm{y}$, is guaranteed according to Thm. 2 and inside $\textbf{B}_{\rho_{-}}$ there exists a region $\bm{\Omega}_{\rho_{-}}$, inside of which positive energy growth can be observed. Then the long-time behavior of solutions to the system (23) with initial conditions $\bm{P}^{\frac{1}{2}}\bm{y}_{0}=\bm{z}_{0}\in\textbf{B}_{\rho_{+}}$ can alternate between $\textbf{B}_{\rho_{-}}\setminus\bm{\Omega}_{\rho_{-}}$ and $\bm{\Omega}_{\rho_{-}}$, as shown in the Fig. 3. In a special case where $\bm{z}_{0}$ is a real solution to Eq. (31a), the long-time behavior of the system stays at the boundary of $\bm{\Omega}_{\rho_{-}}$. It is not difficult to conclude that
@@
which correctly reproduces the original trapping theorem guaranteeing global boundedness (in fact, a monotonically trapping region). We can also validate the limit by setting $\epsilon_{Q}$ to zero so that $\bm{\tilde{}}{\bm{H}}=0$, then the time derivative of the Lyapunov function (31a) becomes
@@
Figure 2: Sketch of the trapping region. $\dot{v}<0$ is observed in $\bm{\Omega}_{\rho_{+}}\setminus\bm{\Omega}_{\rho_{-}}$ and long-term stability is guaranteed for any initial conditions $\|\bm{y}(t_{0})\|_{\bm{P}}=\|\bm{z}(t_{0})\|_{2}\leq\rho_{+}$.
@@
Note that the “shell” of stability $\textbf{B}_{\rho_{+}}\setminus\textbf{B}_{\rho_{-}}$ vanishes when $\rho_{+}=\rho_{-}$ or equivalently when the square root vanishes:
@@ III Extended trapping SINDy algorithm
In this section, we develop the extended trapping SINDy algorithm building on Thm. 2, which enforces local stability guarantees during data-driven model identification. Unlike intrusive methods like POD-Galerkin, which require costly field-derivative evaluations, our approach directly constructs stable models from data via sparse optimization [32, 11].
In particular, we use the trapping SINDy algorithm [2] and formulate a modified “extended” trapping SINDy algorithm. We first review the standard SINDy framework [28] and the trapping SINDy algorithm, which successfully incorporates the trapping theorem (Thm. 1) into model identification in Sec. III.1. We then introduce our proposed extended trapping SINDy algorithm in Sec III.2.
@@ III.1 Standard and trapping SINDy algorithm
where $\dot{\bm{X}}_{:,j}$ denotes the $j^{th}$ column of $\dot{\bm{X}}$, $\|\cdot\|_{0}$ counts the number of nonzero elements and $\gamma$ determines the strength of sparsity-promotion during identification of the $\bm{\Xi}$. One may also notice that $\|\cdot\|_{0}$ is not convex, which leads to a nonconvex optimization. One way to ameliorate this is to instead use the $\ell_{1}$ norm,resulting in the LASSO [39]. The $\ell_{1}$ is convex but promotes a weaker form of sparsity. A computationally efficient alternative to directly solving the $\ell_{0}$-regularized problem was demonstrated in Brunton et al. [28], where a sequential threshold least-squares algorithm was used to greedily solve the problem; its convergence properties were proven in [40].
To encode the enforced energy-preserving constraint on the quadratics terms, Loiseau et al. [41] vectorizes $\bm{\Xi}[:]=\bm{\xi}\in\mathbb{R}^{rN}$ to recast the constrained minimization problem as an unconstrained convex problem using an augmented functional formulation. If the constraint is encoded as $\bm{C}\bm{\xi}=\bm{q}$, together with the improvements above, the optimization problem reads,
@@
where $\bm{\hat{\Theta}}(\bm{X})=\bm{I}\otimes\bm{\Theta}(\bm{X})$, $\dot{\bm{X}}[:]$ is the vectorized $\dot{\bm{X}}$ so that we have $\bm{\hat{\Theta}}\in\mathbb{R}^{rM\times rN}$, $\dot{\bm{X}}[:]\in\mathbb{R}^{rM}$. The number of constraints of the energy-preserving structure is given by $p=r(r+1)(r+2)/6$ and therefore we have $\bm{C}\in\mathbb{R}^{p\times rN}$, $\bm{q}\in\mathbb{R}^{p}$ and the Lagrange multiplier $\bm{\zeta}\in\mathbb{R}^{p}$. Additionally, for the linear equality constraints of the energy-preserving structure, $\bm{q}=0$ and the number of free parameters is given by rNp=2prN-p=2p. If needed, these constraints can be used to impose any additional linear relationships from physical knowledge, such as enforcing global invariant preservation or enforcing a priori known values of any entries $\xi_{i}$.
@@
$\bm{\xi},\bm{\zeta},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}^{\prime}}$ $\max_{\bm{\zeta}}\Biggr[\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (37)
@@
where the new term denotes that the largest eigenvalue of $\tilde{\bm{A}}^{S}$ will be enforced to be negative so that $\tilde{\bm{A}}^{S}$ is Hurwitz, and the magnitude of whose loss is modulated by the hyperparameter $\eta$. One should notice that the last term of this problem is not convex but convex composite, making the problem non-convex, however the details of the algorithm used for solving the optimization problem in (37) will not be introduced in this work; see Kaptanoglu et al. [2] for more details.